OpenStudy (anonymous):

What is the product of the square root of t. (8the square root of 2. - 7)? Simplify if possible.

zepdrix (zepdrix):

So this is what we need to simplify? \[\Large \sqrt t \left(8\sqrt2-7\right)\]

OpenStudy (anonymous):

yes!

zepdrix (zepdrix):

So we can start by `distributing` the sqrt(t) to each term inside the brackets, giving us,\[\Large \left(8\sqrt2\sqrt t-7\sqrt t\right)\]We can drop the brackets since nothing is being applied to the outside of them anymore,\[\Large 8\sqrt2\sqrt t-7\sqrt t\]We can combine the square roots in the first term,\[\Large 8\sqrt{2t}-7\sqrt t\] Hmm we can't really do much simplification beyond that.. :\

OpenStudy (anonymous):

thats all yoou had to do??

OpenStudy (anonymous):

2t the square root of the quantity 2 times t. 8the square root of the quantity 2 times t. - 7 8the square root of the quantity 2 times t. - 7the square root of t.

OpenStudy (anonymous):

those were the answeers

zepdrix (zepdrix):

Hmm does our answer match any of those descriptions? :)

OpenStudy (anonymous):

yeah! the last one (:

zepdrix (zepdrix):

Yay good job. The only really tricky part is distributing the sqrt(t). Make sure you're comfortable with that step.

OpenStudy (anonymous):

yeah i need more practice with that

OpenStudy (preetha):

Tom, dont forget to give Zedprix a medal!